# BRcal: An R Package to Boldness-Recalibrate Probability Predictions

### v2

@@ A.1.1 Nested LLO-Adjustment
$=c(x_{i};\delta^{\prime}\delta^{\gamma^{\prime}},\gamma,\gamma^{\prime})$
@@ A.1.3 Log LLO-Adjustment
The following results pertains to log LLO-adjusted predictions needed for calculations of the gradients in Sections A.2 and A.3.
@@ A.1.4 Intermediate Partial Derivatives
The following results pertain to derivatives needed for calculations of the gradients in Sections A.2 and A.3.
@@
A.2 Objective Function Gradients
The following series of equations show the derivation of the gradients of the objective function in Equation 10 using the results from Section A.1.
@@ A.2 Objective Function Gradients
A.3 Constraint Function Gradients
The following series of equations show the derivation of the gradients of the constraint function in Equation 12 using the results from Section A.1.
$\frac{\partial f(\delta,\gamma)}{\partial\delta}$ $=\frac{1}{\left(1+BF\frac{P(M_{u})}{P(M_{c})}\right)^{2}}\frac{\partial}{\partial\delta}\left(1+BF\frac{P(M_{u})}{P(M_{c})}\right)$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})^{2}\frac{P(M_{u})}{P(M_{c})}exp\left\{-\frac{1}{2}(BIC_{u}-BIC_{c})\right\}\frac{\partial}{\partial\delta}-\frac{1}{2}(BIC_{u}-BIC_{c})$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})^{2}\left(\frac{1}{P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})}-1\right)$
@@ A.3 Constraint Function Gradients
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))$
@@
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\left(\sum_{i=1}^{n}\frac{\hat{\gamma}}{\delta}y_{i}-\frac{\hat{\gamma}}{\delta}x_{i,MLE}^{\prime}-\frac{1}{\delta}y_{i}+\frac{1}{\delta}x_{i}^{\prime}\right)$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\frac{1}{\delta}\left(\sum_{i=1}^{n}(\hat{\gamma}-1)y_{i}-\hat{\gamma}x_{i,MLE}^{\prime}+x_{i}^{\prime}\right)$
$\frac{\partial f(\delta,\gamma)}{\partial\gamma}$ $=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))$
@@
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}\hat{\gamma}y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)(1-x_{i,MLE}^{\prime})$
@@
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}\hat{\gamma}y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)-\hat{\gamma}log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i,MLE}^{\prime}$
@@
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)\big{(}(\hat{\gamma}-1)y_{i}-\hat{\gamma}x_{i,MLE}^{\prime}+x_{i}^{\prime}\big{)}\bigg{)}$
@@ Appendix B Strategy for Specifying Parameter Grid in plot_params()
When using plot_params(), some care is required for the specifications of arguments dlim, glim, and k, which together determine the fineness of your grid of parameter values plotted. We suggest the following scheme for determining the grid of parameter values to search. First, fix k at some small number, less than 20 for sake of computation time. Then, center a grid with small range around $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$ for $\mathbf{x}$. Increase the size of k until your grid detects approximated the probability of calibration at the MLEs that you expect. At this point, you may find that your grid does not fully encompass the range of parameters values which produce recalibrated probabilities with meaningfully high posterior probability of being calibrated. If this is the case, expand your grid until it the region with high probability of calibration is covered. Then, increase k to create a fine grid of values. Additionally, we caution users from including $\gamma=0$ in the grid. This setting recalibrates all values in $\mathbf{x}$ to a single value which is not desirable in practice. Unless the singular value is near the base rate, the set will be poorly calibrated and minimally bold, which does not align with the goal of boldness-recalibration.
As mentioned, the computation time to construct this plot can be high if the grid of parameter values is very fine or the sample size of predictions is very large. In an effort to reduce this bottleneck, we have developed as strategy to circumvent unnecessary calls to the optim(). To calculate $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$ for each grid cell, this requires evaluating the likelihood at the MLEs for each LLO-adjusted set $\mathbf{x}^{\prime}_{ij}$, which in turn requires a call to optim() to get the MLEs. However, given the MLEs for the original set $\mathbf{x}$ and the $\delta_{i}$ and $\gamma_{j}$ from the grid, we show that it is possible to get the MLEs for $\mathbf{x}^{\prime}_{ij}$ without optimization via the result in Appendix C. Thus, we can reduce the number of optim() calls from one call for each grid cell to just one call overall to find the MLEs for the original set. An example of a direct result of this is reducing the computation time over a grid for n=868, k=100 from approximately 2 hours to under a minute.
@@
$P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$ $=\left(1+\frac{P(M_{u})}{P(M_{c})}exp\{-\frac{1}{2}(BIC_{u}-BIC_{c})\}\right)^{-1}$
where $BIC_{u}=2\times log(n)-2\times log(\pi(\hat{\delta}_{MLE},\hat{\gamma}_{MLE}|\mathbf{x}_{ij}^{\prime},\mathbf{y})))$, $\mathbf{x}_{ij}^{\prime}=c(\mathbf{x};\delta_{i},\gamma_{j})$, and $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ are the MLEs for $\mathbf{x}_{ij}^{\prime}$. Notice that $BIC_{u}$ requires a unique calculation of $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ for each $i,j$. In this section, we show that $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ can be calculated using only the MLEs for $\mathbf{x}$ ($\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$), and the grid shift parameters ($\delta_{i}$ and $\gamma_{j}$). This is important as it circumvents an additional call to optim() for each grid cell.

### v3

@@ A.1.1 Nested LLO-Adjustment
$=c(x_{i};\delta^{\prime}\delta^{\gamma^{\prime}},\gamma\gamma^{\prime})$
@@ A.1.3 Log LLO-Adjustment
The following results pertains to log LLO-adjusted predictions needed for calculations of the gradients in Sections A.3 and A.4.
@@ A.1.4 Intermediate Partial Derivatives
The following results pertain to derivatives needed for calculations of the gradients in Sections A.3 and A.4.
@@
A.2 MLE Objective Function Gradients
The following series of equations show the derivation of the gradients of the objective function in (42) using the results from Section A.1. We will derive these in terms of $\delta$ and then reparameterize in terms of $\tau$.
$\frac{\partial h(\delta,\gamma)}{\partial\delta}$ $=-\sum_{i=1}^{n}y_{i}\frac{\partial}{\partial\delta}log(x_{i}^{\prime})+(1-y_{i})\frac{\partial}{\partial\gamma}log(1-x_{i}^{\prime})$
$=-\sum_{i=1}^{n}y_{i}\frac{1}{\delta}(1-x_{i}^{\prime})+(1-y_{i})(-\frac{1}{\delta}x_{i}^{\prime})$
$=\sum_{i=1}^{n}\frac{1}{\delta}(x_{i}^{\prime}(1-y_{i})-y_{i}(1-x_{i}^{\prime}))$
$=\sum_{i=1}^{n}\frac{1}{\delta}(x_{i}^{\prime}-y_{i})$
$=\sum_{i=1}^{n}\frac{1}{\delta}(c(x_{i};\delta,\gamma)-y_{i})$
$\implies\frac{\partial h(\tau,\gamma)}{\partial\tau}$ $=\sum_{i=1}^{n}\frac{1}{exp(\tau)}(c(x_{i};exp(\tau),\gamma)-y_{i})\frac{\partial}{\partial\tau}exp(\tau)$
$=\sum_{i=1}^{n}(c(x_{i};exp(\tau),\gamma)-y_{i})$
$\frac{\partial h(\delta,\gamma)}{\partial\gamma}$ $=-\sum_{i=1}^{n}y_{i}\frac{\partial}{\partial\gamma}log(x_{i}^{\prime})+(1-y_{i})\frac{\partial}{\partial\gamma}log(1-x_{i}^{\prime})$
$=-\sum_{i=1}^{n}y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)(1-x_{i}^{\prime})+(1-y_{i})(-log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i}^{\prime})$
$=\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)(x_{i}^{\prime}(1-y_{i})-y_{i}(1-x_{i}^{\prime}))$
$=\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)(x_{i}^{\prime}-y_{i})$
$=\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)(c(x_{i};\delta,\gamma)-y_{i})$
$\implies\frac{\partial h(\tau,\gamma)}{\partial\gamma}$ $=\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)(c(x_{i};exp(\tau),\gamma)-y_{i})$
A.3 Boldness-Recalibration Objective Function Gradients
The following series of equations show the derivation of the gradients of the objective function in (11) using the results from Section A.1.
@@ A.3 Boldness-Recalibration Objective Function Gradients
Thus, we define the Jacobian of (11) as $J_{f}(\delta,\gamma)=\left[\frac{\partial f}{\partial\delta},\frac{\partial f}{\partial\gamma}\right]$ where
$\frac{\partial f}{\partial\gamma}$ $=\frac{-1}{(n-1)sd(\mathbf{x}^{\prime})}\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\left(log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i}^{\prime}(1-x_{i}^{\prime})-\frac{1}{n}\sum_{j=1}^{n}log\left(\frac{x_{j}}{1-x_{j}}\right)x_{j}^{\prime}(1-x_{j}^{\prime})\right).$ (15)
A.4 Boldness-Recalibration Constraint Function Gradients
The following series of equations show the derivation of the gradients of the constraint function in (13) using the results from Section A.1.
$\frac{\partial g(\delta,\gamma)}{\partial\delta}$ $=\frac{1}{\left(1+BF\frac{P(M_{u})}{P(M_{c})}\right)^{2}}\frac{\partial}{\partial\delta}\left(1+BF\frac{P(M_{u})}{P(M_{c})}\right)$
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})^{2}\frac{P(M_{u})}{P(M_{c})}exp\left\{-\frac{1}{2}(BIC_{u}-BIC_{c})\right\}\frac{\partial}{\partial\delta}-\frac{1}{2}(BIC_{u}-BIC_{c})$
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})^{2}\left(\frac{1}{P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})}-1\right)$
@@ A.4 Boldness-Recalibration Constraint Function Gradients
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))$
@@
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))\left(\sum_{i=1}^{n}\frac{\hat{\gamma}}{\delta}y_{i}-\frac{\hat{\gamma}}{\delta}x_{i,MLE}^{\prime}-\frac{1}{\delta}y_{i}+\frac{1}{\delta}x_{i}^{\prime}\right)$
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))\frac{1}{\delta}\left(\sum_{i=1}^{n}(\hat{\gamma}-1)y_{i}-\hat{\gamma}x_{i,MLE}^{\prime}+x_{i}^{\prime}\right)$
$\frac{\partial g(\delta,\gamma)}{\partial\gamma}$ $=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))$
@@
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))\bigg{(}\sum_{i=1}^{n}\hat{\gamma}y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)(1-x_{i,MLE}^{\prime})$
@@
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))\bigg{(}\sum_{i=1}^{n}\hat{\gamma}y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)-\hat{\gamma}log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i,MLE}^{\prime}$
@@
$=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))\bigg{(}\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)\big{(}(\hat{\gamma}-1)y_{i}-\hat{\gamma}x_{i,MLE}^{\prime}+x_{i}^{\prime}\big{)}\bigg{)}$
Thus, we define the Jacobian of (13) as $J_{g}(\delta,\gamma)=\left[\frac{\partial g}{\partial\delta},\frac{\partial g}{\partial\gamma}\right]$ where
$\frac{\partial g}{\partial\delta}$ $=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))\frac{1}{\delta}\left(\sum_{i=1}^{n}\hat{\gamma}_{MLE}^{\prime}(y_{i}-x_{i,MLE}^{\prime})-y_{i}+x_{i}^{\prime}\right),$ (16)
$\frac{\partial g}{\partial\gamma}$ $=P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})(1-P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}))\bigg{(}\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)\big{(}\hat{\gamma}_{MLE}^{\prime}(y_{i}-x_{i,MLE}^{\prime})-y_{i}+x_{i}^{\prime}\big{)}\bigg{)}.$ (17)
@@ Appendix B Strategy for Specifying Parameter Grid in plot_params()
When using plot_params(), some care is required for the specifications of arguments dlim, glim, and k, which together determine the fineness of your grid of parameter values plotted. We suggest the following scheme for determining the grid of parameter values to search. First, fix k at some small number, less than 20 for sake of computation time. Then, center a grid with small range around $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$ for $\mathbf{x}$. A grid that is too course may not be able to detect regions of non-zero probability of calibration. Increase the grid fineness (i.e., increase k) until the probability of calibration near the MLEs on the plot approximately matches the probability of calibration from bayes_ms() for the same set of data. At this point, you may find that your grid does not fully encompass the range of parameters values which produce recalibrated probabilities with meaningfully high posterior probability of being calibrated. If this is the case, expand your grid until it the region with high probability of calibration is covered. Then, further increase k to create a fine grid of values. Additionally, we caution users from including $\gamma=0$ in the grid. This setting recalibrates all values in $\mathbf{x}$ to a single value which is not desirable in practice. Unless the singular value is near the base rate, the set will be poorly calibrated and minimally bold, which does not align with the goal of boldness-recalibration.
As mentioned, the computation time to construct this plot can be high if the grid of parameter values is very fine or the sample size of predictions is very large. In an effort to reduce this bottleneck, we have developed a strategy to circumvent unnecessary calls to optim(). To calculate $P_{\mathbf{x}^{\prime}_{ij}}(M_{c}|\mathbf{y})$ for each grid cell, this requires evaluating the likelihood at the MLEs for each LLO-adjusted set $\mathbf{x}^{\prime}_{ij}$, which in turn requires a call to optim() to get the MLEs. However, given the MLEs for the original set $\mathbf{x}$ and the $\delta_{i}$ and $\gamma_{j}$ from the grid, we show that it is possible to get the MLEs for $\mathbf{x}^{\prime}_{ij}$ without optimization via the result in Appendix C. Thus, we can reduce the number of optim() calls from one call for each grid cell to just one call overall to find the MLEs for the original set. A direct result of this is that this reduces the computation time from 12.03 minutes to 32.88 second for the hockey data and from 1.03 hours to 2.26 minutes for the foreclosure data, on average for k=200 (on Ubuntu 22.04, R version 4.4.2).
@@
$P_{\mathbf{x}^{\prime}_{ij}}(M_{c}|\mathbf{y})$ $=\left(1+\frac{P(M_{u})}{P(M_{c})}exp\{-\frac{1}{2}(BIC_{u}-BIC_{c})\}\right)^{-1}$
where $BIC_{u}=2\times log(n)-2\times log(\pi(\hat{\delta}^{\prime}_{MLE},\hat{\gamma}^{\prime}_{MLE}|\mathbf{x}_{ij}^{\prime},\mathbf{y})))$, $\mathbf{x}_{ij}^{\prime}=c(\mathbf{x};\delta_{i},\gamma_{j})$, and $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ are the MLEs for $\mathbf{x}_{ij}^{\prime}$. Notice that $BIC_{u}$ requires a unique calculation of $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ for each $i,j$. In this section, we show that $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ can be calculated using only the MLEs for $\mathbf{x}$ ($\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$), and the grid shift parameters ($\delta_{i}$ and $\gamma_{j}$). This is important as it circumvents an additional call to optim() for each grid cell.